# Theoretical study of transition energies and matrix elements of the cadmium atom

### v1

@@ Abstract
[…]
@@ I Introduction
[…]
@@ II.1 Fock-space coupled cluster
Here $D_{K}$ is the initial energy denominator, $\mathnormal{S_{K}}$ is a shift parameter (in the present case, we resorted to real shifts [28]) and $\mathnormal{n}$ is a non-negative integer parameter. Note that $\mathnormal{S_{K}}$ should be chosen to be relatively small compared to large $\mathnormal{D_{K}}$ values, while still being substantial enough to ensure that the shifted denominators remain negative and sufficiently far from zero to avoid divergences. A very large active space was considered in order to obtain the highly lying excited states, including the $n=(5-9)s$, $n=(5-8)d$, and $n=(5-16)p$ particle orbitals, accounting for 110 active bispinors. The minimum energy-shift parameters that allowed for convergence within this active space were --0.1 and --0.25 for the $0h1p$ and $0h2p$ sectors, respectively. Previously, the denominator shifting technique was successfully developed [29, 30, 31, 32] and employed [33] to avoid divergences for electronic structure calculations in $0h2p$ Fock space sector. For a more detailed and comprehensive description of the technique, one may refer to Ref. [7].
To refine the results of the computation, we also performed calculation within a framework of the Intermediate-Hamiltonian (IH) Fock-space coupled cluster method for incomplete main model spaces [34]. This approach involves partitioning the model space into a main model space and an intermediate model space, such that the contributions of the intermediate space functions to the target states are minimal, thereby avoiding the intruder-state problem. For more details, see Ref. [34]. In this calculation, the total model space consisted of 224 lowest-lying particle orbitals, while the main model space was chosen to comprise determinants with configurations $n=(5-7)s$, $n=(5-7)d$, and $n=(5-14)p$ in sector $0h1p$ and configurations $5sn$, where $n=(5-7),(5-7)d,(5-14)p$, in sector $0h2p$.
@@ III Results and discussion
Table 1: Electronic transition energies for the considered states relative to the ground state (in cm-1). Column labeled ‘‘FS-CCSD (shifts)’’ shows the results obtained using the FS-CC method with regular energy-denominator shifts, column labeled "FS-CCSD (IH)" shows the results obtained using the Intermediate-Hamiltonian FS-CC method.
@@
As seen in Table 2, the present calculations agree well with the available theoretical and experimental data for the low-lying transitions. For the $5s^{2}\>{}^{1}S_{0}\to 5s5p\>^{3}P^{o}_{1}$ transition, both FS-CCSD approaches give values closer to the experimental one. In the case of $5s^{2}\>{}^{1}S_{0}\to 5s5p\>^{1}P^{o}_{1}$ transition, all theoretical results are quite close, while having a distinct gap with the available experimental value. In contrast, the $5s^{2}\>{}^{1}S_{0}\to 5s6p\>^{1}P^{o}_{1}$ transition values given by both FS-CCSD methods are noticeably far from the very close values given by different versions of the CI method. The largest discrepancy for this transition is found between the IH FS-CCSD and DFCP [19] methods, and amounts to about 10%. For the $5s6s\>^{1}S_{0}\to 5s5p\>^{1}P^{o}_{1}$ transition, the present CI calculation agrees well with the result from Ref. [21]. The difference with the corresponding FS-CCSD calculations is about 5%. The values obtained for the $5s5d\>^{1}D_{2}\to 5s5p\>^{1}P^{o}_{1}$ transition closely match other available theoretical results, with the FS-CCSD calculations being closer to the isolated experimental value.
@@
$5s10p\>^{1}P^{o}_{1}$ 0.008 0.072 0.085
If we analyze the results obtained for the transitions involving higher-lying states from Table 3, we can see that, in general, two different coupled cluster approaches agree to a great extent for transitions from all 3 considered states ($5s^{2}\>{}^{1}S$, $5s6s\>^{1}S$ and $5s5d\>^{1}D$) to the states up to $5s8p\>^{3}P^{o}_{1}$. In the case of transitions from the $5s^{2}\>{}^{1}S_{0}$ state, the correspondence between the values for singlet $\to$ singlet transitions given by the FS-CC method with energy-denominator shifts (denoted as FS-CC (shifts)) and CI method is good up to $5s8p\>^{1}P^{o}_{1}$. For the IH FS-CC method, the agreement with CI extends up to the $5s11p:^{1}P^{o}_{1}$ state. However, for the forbidden singlet $\to$ triplet transitions the situation is much less satisfactory: even for the $5s^{2}\>{}^{1}S_{0}\to 5s6p\>^{3}P^{o}_{1}$ transition the values obtained by the coupled-cluster approaches differ by a factor of two compared to the CI method. Since the transition matrix elements for this transition are already quite small, we suggest that more thorough computations are needed for all higher singlet $\to$ triplet transitions. For the $5s6s\>^{1}S_{0}\to 5snp\>^{1}P^{o}_{1}$ transitions, good agreement is found up to $n=11$ when comparing IH FS-CCSD and CI+MBPT methods, and up to $n=8$ for FS-CCSD (shifts) and CI+MBPT methods. In contrast, for the $5s6s\>^{1}S_{0}\to 5snp\>^{3}P^{o}_{1}$ transitions, the two methods yield noticeably different results for $n>8$. Good agreement between the IH FS-CC and CI methods for singlet $\to$ singlet transitions is also observed for the transitions up to $5s5d\>^{1}D_{2}\to 5s10p\>^{1}P^{o}_{1}$, although in the case of singlet $\to$ triplet transitions, the discrepancy of less than 10% is achieved only for transitions up to the $5s5d\>^{1}D_{2}\to 5s8p\>^{3}P^{o}_{1}$ transition, with a more noticeable difference still present for the $5s5d\>^{1}D_{2}\to 5s5p\>^{3}P^{o}_{1}$.
Comparing the performance of the FS-CC (shifts) and IH FS-CC methods, it can be seen from Table 2 that a relatively good agreement between these implementations for singlet $\to$ singlet transitions is achieved up to the $5s^{2}/5s6s\>^{1}S_{0}\to 5s9p\>^{1}P^{o}_{1}$ transition for those involving the $5s^{2}\>{}^{1}S_{0}$ and $5s6s\>^{1}S_{0}$ states and up to the $5s5d\>^{2}D_{2}\to 5s8p\>^{1}P^{o}_{1}$ for those involving the $5s^{2}\>{}^{1}S_{0}$ state. In these cases, the IH FS-CC method yields results that are generally closer to the CI+MBPT values, particularly for higher-lying transitions. For singlet $\to$ triplet transition, except for the transitions involving low-lying states up to the $5s8p\>^{3}P^{o}_{1}$, all three methods give noticeably different results, with the values computed by IH FS-CC generally showing better agreement with those obtained from CI+MBPT.
@@
Using the matrix elements calculated in the present work and the same laser frequencies ($\omega_{1}=\omega_{2}=375$ nm, $\omega_{3}=710$ nm), along with the same parameters for the shape factor $S(\omega_{1}+\omega_{2})$ for the two FWM schemes (using $5s6s\>^{1}S_{0}$ or $5s5d\>^{2}D_{2}$ as an intermediate state), considered in Ref. [1] for the spectroscopy of the 229Th isomer, the third-order nonlinear susceptibility $\chi_{a}^{(3)}$ of Cd vapor can be computed. The results obtained in the case of two-photon resonance for the pure 114Cd isotope are presented in Table 4.
Table 4: Absolute values of the third-order nonlinear susceptibility of Cd vapor $|\chi_{a}^{(3)}|$ in units of $10^{-6}(ea_{0})^{4}$cm3 and the corresponding values of factors $|\chi_{12}|$ in units of $10^{-15}(ea_{0})^{2}$z-1 and $|\chi_{34}|$ in units of $10^{-16}(ea_{0})^{2}$z-1 for the pure 114Cd utilizing two FWM schemes with $5s6s\>^{1}S_{0}$ and $5s5d\>^{2}D_{2}$ as intermediate states.
@@ IV Conclusions
The values of the third-order nonlinear susceptibility obtained within the IH FS-CCSD and CI+MBPT methods agree within 5% for both considered FWM schemes. As a reference for the $|\chi^{(3)}_{a}|$, we suggest using the IH FS-CCSD results with an estimated uncertainty of 5%. These values are expected to be of significant interest for the preparation of neutral Cd spectroscopy experiment, as well as for the generation of laser light at 148.4 nm through four-wave mixing in cadmium vapor, aimed at high-precision spectroscopy of the 229Th isomeric transition. The discrepancy between the FS-CC and CI results for high-lying transitions highlights the need for more advanced experimental and theoretical approaches to accurately determine considered properties.
@@ Acknowledgments
We thank Qi Xiao and Mikhail Reiter for discussions and comments. Penyazkov and Ding are supported by the National Natural Science Foundation of China (No. 12341401 and No. 12274253) and the Shanghai Municipal Science and Technology Commission (No. 25LZ2600402), Yu is supported by the National Key Research and Development Program of China (2021YFA1402104) and a Project funded by the Space Application System of China Manned Space Program. Skripnikov is supported by Russian Science Foundation under grant  No. 24-12-00092, https://rscf.ru/project/24-12-00092/.

### v2

@@ Abstract
The relativistic Fock-space coupled-cluster methods are applied to the cadmium atom. A large number of transition energies and matrix elements are calculated for the $5s^{2}\>$$^{1}S\to 5snp\>^{1,3}P^{o}$, $5s6s\>^{1}S\to 5snp\>^{1,3}P^{o}$ and $5s5d\>^{1}D\to 5snp\>^{1,3}P^{o}$ transitions for a wide range of $p$ states accounting for relativistic and electron-correlation effects. The results obtained within two different approaches (Fock-space coupled cluster and configuration interaction) are compared with available experimental and theoretical data. Good agreement is found between the two methods for transitions involving low-lying excited $p$-states, whereas for high-lying states, the discrepancy becomes large. The calculated values are used to determine the third-order nonlinear susceptibility of cadmium vapor, with an agreement within 5% between the different methods. The results of the present computations support the feasibility of generating vacuum ultraviolet light in Cd vapor via a four-wave mixing process for the direct excitation of 229Th nucleus.
@@ I Introduction
Experimental estimates of transition dipole moments in Cd are only available for only a few transitions [6]. Previously, knowledge of transition rates relied primarily on several theoretical studies, determining dipole moment transition matrix elements in cadmium. In Refs. [15, 16], the configuration interaction method with many-body perturbation theory (CI+MBPT) [17] was used to calculate the transition matrix elements for transitions from the ground $5s^{2}\>$$^{1}S$ and excited $5s5p\>^{3}P^{o}$ states to the lowest-lying spin-allowed states, up to the $5s7s\>^{3}S$ state. In the work [16] these values were also computed within a CI framework, combined with the all-order linearized coupled-cluster method [18]. Later, in Ref. [19], the Dirac-Fock plus core polarization [20] and relativistic configuration interaction methods were employed to obtain transition matrix elements for the same set of transitions. In the recent work [21], the set of calculated transition matrix elements has been expanded within the relativistic CI+MPBT framework to include low-lying transitions from the $5s5p\>^{3}P^{o}_{1}$ and $5s5p\>^{3}P^{o}_{2}$ states. All of the aforementioned calculations use similar methodology and yield consistent results. However, there is still no theoretical or experimental data available for transition matrix elements involving high-lying states in general.
@@ II.1 Fock-space coupled cluster
Here $D_{K}$ is the initial energy denominator, $\mathnormal{S_{K}}$ is a shift parameter (in the present case, we resorted to real shifts [28]) and $\mathnormal{n}$ is a non-negative integer parameter. Note that $\mathnormal{S_{K}}$ should be chosen to be relatively small compared to large $\mathnormal{D_{K}}$ values, while still being substantial enough to ensure that the shifted denominators remain negative and sufficiently far from zero to avoid divergences. A very large active space was considered in order to obtain the highly lying excited states, including the n=(59)sn=(5-9)s, n=(58)dn=(5-8)d, and n=(516)pn=(5-16)p particle orbitals, accounting for 110 active bispinors. The minimum energy-shift parameters that allowed for convergence within this active space were -0.1 and -0.25 for the $0h1p$ and $0h2p$ sectors, respectively. Previously, the denominator shifting technique was successfully developed [29, 30, 31] and employed [32] to avoid divergences for electronic structure calculations in $0h2p$ Fock space sector. For a more detailed and comprehensive description of the technique, one may refer to Ref. [7].
To refine the results of the computation, we also performed calculation within a framework of the Intermediate-Hamiltonian (IH) Fock-space coupled cluster method for incomplete main model spaces [31]. This approach involves partitioning the model space into a main model space and an intermediate model space, such that the contributions of the intermediate space functions to the target states are minimal, thereby avoiding the intruder-state problem. For more details, see Ref. [31]. In this calculation, the total model space consisted of 224 lowest-lying particle orbitals, while the main active space was chosen to comprise determinants with configurations n=(57)sn=(5-7)s, n=(57)dn=(5-7)d, and n=(514)pn=(5-14)p in sector $0h1p$ and configurations $5sn$, where n=(57),(57)d,(514)pn=(5-7),(5-7)d,(5-14)p, in sector $0h2p$.
@@ III Results and discussion
Table 1: Electronic transition energies for the considered states relative to the ground state (in cm-1). Column labeled ‘‘FS-CCSD (shifts)’’ shows the results obtained using the FS-CC method with regular energy-denominator shifts, column labeled "FS-CCSD (IH)" shows the results obtained using the Intermediate-Hamiltonian FS-$C$ method. The ’’Basis set’’ column represents a correction to the transition energy accounting for more extensive basis set. For details see Section II.
@@
[…]
@@
If we analyse the results obtained for the transitions involving higher-lying states from Table 3, we can see that, in general, two different coupled cluster approaches agree within $\sim 15\%$ for transitions from all 3 considered states ($5s^{2}\>{}^{1}S$, $5s6s\>^{1}S$ and $5s5d\>^{1}D$) to the states up to $5s8p\>^{3}P^{o}_{1}$. In the case of transitions from the $5s^{2}\>{}^{1}S_{0}$ state, the correspondence between the values for singlet $\to$ singlet transitions given by the FS-CC method with energy-denominator shifts (denoted as FS-CC (shifts)) and CI method is less than 15% up to $5s8p\>^{1}P^{o}_{1}$. For the IH FS-CC method, the agreement with a deviation of no more than 15% with CI extends up to the $5s11p\>^{1}P^{o}_{1}$ state. However, for the forbidden singlet $\to$ triplet transitions the situation is much less satisfactory: even for the $5s^{2}\>{}^{1}S_{0}\to 5s6p\>^{3}P^{o}_{1}$ transition the values obtained by the coupled-cluster approaches differ by a factor of two compared to the CI method. Since the transition matrix elements for this transition are already quite small, we suggest that more thorough computations are needed for all higher singlet $\to$ triplet transitions. For the $5s6s\>^{1}S_{0}\to 5snp\>^{1}P^{o}_{1}$ transitions, reasonable agreement is found up to $n=11$ when comparing IH FS-CCSD and CI+MBPT methods, and up to $n=8$ for FS-CCSD (shifts) and CI+MBPT methods. In contrast, for the $5s6s\>^{1}S_{0}\to 5snp\>^{3}P^{o}_{1}$ transitions, the two methods yield results which differ by at least a factor of 2 for $n>8$. A 15% agreement between the IH FS-CC and CI methods for singlet $\to$ singlet transitions is also observed for the transitions up to $5s5d\>^{1}D_{2}\to 5s10p\>^{1}P^{o}_{1}$, although in the case of singlet $\to$ triplet transitions, the discrepancy of less than 10% is achieved only for transitions up to the $5s5d\>^{1}D_{2}\to 5s8p\>^{3}P^{o}_{1}$ transition, with a difference of about 25% still present for the $5s5d\>^{1}D_{2}\to 5s5p\>^{3}P^{o}_{1}$ transition.
Comparing the performance of the FS-CC (shifts) and IH FS-CC methods, it can be seen from Table 2 that a relatively good agreement between these implementations for singlet $\to$ singlet transitions is achieved up to the $5s^{2}/5s6s\>^{1}S_{0}\to 5s9p\>^{1}P^{o}_{1}$ transition for those involving the $5s^{2}\>{}^{1}S_{0}$ and $5s6s\>^{1}S_{0}$ states and up to the $5s5d\>^{2}D_{2}\to 5s8p\>^{1}P^{o}_{1}$ for those involving the $5s^{2}\>{}^{1}S_{0}$ state. In these cases, the IH FS-CC method yields results that are generally closer to the CI+MBPT values, particularly for higher-lying transitions. For singlet $\to$ triplet transition, except for the transitions involving low-lying states up to the $5s8p\>^{3}P^{o}_{1}$, all three methods give noticeably different results, with the values computed by IH FS-CC generally showing better agreement with those obtained from CI+MBPT. All considered methods agree with the experiment within the experimental uncertainty.
@@
Using the matrix elements calculated in the present work and the same laser frequencies ($\omega_{1}=\omega_{2}=375$ nm, $\omega_{3}=710$ nm), along with the same parameters for the shape factor $S(\omega_{1}+\omega_{2})$ for the two FWM schemes (using $5s6s\>^{1}S_{0}$ or $5s5d\>^{2}D_{2}$ as an intermediate state), considered in Ref. [1] for the direct excitation of the 229Th nuclear transition, the third-order nonlinear susceptibility $\chi_{a}^{(3)}$ of Cd vapor can be computed. The results obtained in the case of two-photon resonance for the pure 114Cd isotope are presented in Table 4.
IV Conclusions
Table 4: Absolute values of the third-order nonlinear susceptibility of Cd vapor $|\chi_{a}^{(3)}|$ in units of $10^{-6}(ea_{0})^{4}$cm3 and the corresponding values of factors $|\chi_{12}|$ in units of $10^{-15}(ea_{0})^{2}$Hz-1 and $|\chi_{34}|$ in units of $10^{-16}(ea_{0})^{2}$Hz-1 for the pure 114Cd utilizing two FWM schemes with $5s6s\>^{1}S_{0}$ and $5s5d\>^{2}D_{2}$ as intermediate states.
@@ IV Conclusions
[…]
@@ Acknowledgments
We thank Qi Xiao and Mikhail Reiter for discussions and comments. Penyazkov and Ding are supported by the National Natural Science Foundation of China (No. 12341401 and No. 12274253) and the Shanghai Municipal Science and Technology Commission (No. 25LZ2600402), Yu is supported by the National Key Research and Development Program of China (2021YFA1402104), the Innovation Program 659 for Quantum Science and Technology (Grant No. 660 2021ZD0300901), and a Project funded by the Space Application System of China Manned Space Program. Skripnikov is supported by Russian Science Foundation under grant  No. 24-12-00092, https://rscf.ru/project/24-12-00092/.